One mole of an ideal gas undergoes a reversible ad
Answer #1

Question 1

In case of reversible adiabatic we have the formula as

TVγ-1 = constant

So we have

T1V1γ-1=T2V2γ-1

So T1/T2 = (V2/V1)γ-1

T1/T2 = 3γ-1 , So we know that value of of 0<γ-1<1 , SO 3γ-1 >1 , So T1 > T2.

Hence option C is correct